A discount store sold plastic cups for $3.50 each and ceramic cups for $4 each. If 400 cups were sold for $1458, how many cups of each type were sold?
Answer by richwmiller(17219)

p=plastic cups
and c=ceramic
p+c=400
c=400-p
3.50p+4*c=1458
3.50p+4*(400-p)=1458
3.5p+1600-4p=1458
-.5p=-142
p=284
c=116